# Why Go Full? Elevating Federated Learning Through Partial Network Updates

### v2

@@ F.1 Learning Rate Tuning
From the results, it can be seen that both FNU and PNU methods perform best with a learning rate of 0.001. So in our experimental configurations, the Adam optimizer with a learning rate of 0.001 is finally chosen.
@@ Appendix G Justification of Assumption 3
We begin with analysing the lower bound of $k$. Since $S_{1}$ and $S_{2}$ are arbitrary, it possible that $S_{1}=S_{2}$, indicating a lower bound of 1 for the value $k$. As for approximating the upper bound of the $k$, we conduct Monte Carlo simulations on real-world nueral networks.
We test the $k$ values in three neural networks at different training stages. For each neural networks, we conduct Monte Carlo simulations for 10,000 samples to accurately approximate the value of $k$. The experimental results are shown in Table 13. We can see that $k$ is close to 1 under different settings, which proves that the effect of applying different masks to the variability of gradient is similar, thus strongly supporting Assumption 3.

### v3

@@ F.1 Learning Rate Tuning
From the results, it can be seen that both FNU and PNU methods perform best with a learning rate of 0.001. So in our experimental configurations, we ultimately select the Adam optimizer with a learning rate of 0.001.
@@ Appendix G Justification of Assumption 3
We begin with analysing the lower bound of $k$. Since $S_{1}$ and $S_{2}$ are arbitrary, it is possible that $S_{1}=S_{2}$, indicating a lower bound of 1 for the value $k$. As for approximating the upper bound of the $k$, we conduct Monte Carlo simulations on real-world nueral networks.
We test the $k$ values in three neural networks at different training stages. For each neural network, we conduct Monte Carlo simulations to collect 10,000 samples to accurately approximate the value of $k$. The experimental results are shown in Table 13. We can see that $k$ is close to 1 under different settings, which proves that the effect of applying different masks to the variability of gradient is similar, thus strongly supporting Assumption 3.